3. Amaterial has a mass of 95 g and 68 g before and after abrasion, respectively. If the abrasive




ratio of 1.85, find the area of abrasion

1
Expert's answer
2022-02-11T13:28:02-0500

Area of abrasion= ("\frac{185}{100}""\times" 95)-("\frac{185}{100}""\times" 68)


Area of abrasion= 175.75- 125.80


Area of abrasion= 49.95